OpenStudy (anonymous):

Write and factor a binomial that is the difference of 2 perfect squares and that has a greatest common factor of 5mk. I will give out medals!~

OpenStudy (anonymous):

are m and k your variables?

OpenStudy (anonymous):

nd

OpenStudy (anonymous):

Yes*

OpenStudy (anonymous):

Ill give it a whirl, x^2-y^2 for x=5mk and y=5x... just a guess honatsly

OpenStudy (anonymous):

I see and um is x^2-y^2 your final answer?

OpenStudy (anonymous):

I'm a bit confused ^.^;

OpenStudy (anonymous):

(x+y)(x-y)= (5mk+25mk)(5mk-25mk) = (30mk)(-20mk) =-60mk... like i said im just guessing, hope someone else corrects if im wrong, but to me this satisfies the conditions

OpenStudy (anonymous):

Hmm...I guess?

OpenStudy (anonymous):

does it define m and k?

OpenStudy (anonymous):

Supposely if m and k were not separate variables

OpenStudy (anonymous):

Either way, I like your answer, i'm sure you are correct c:

OpenStudy (anonymous):

Thanks anyway~

OpenStudy (anonymous):

you could do m^2-25k^2 where m=5k, np

OpenStudy (anonymous):

good luck :)

OpenStudy (anonymous):

Thanks (: